### v2

# How to Use Large Language Models for Text Coding: The Case of Fatherhood Roles in Public Policy Documents

Abstract
Recent advances in large language models (LLMs) like GPT-3 and GPT-4 promise automation with better results and less programming, opening up new opportunities for text analysis in political science. In this study, we evaluate LLMs on three original coding tasks of non-English political science texts, and we provide a detailed description of a general workflow for using LLMs for text coding in political science research. Our use case offers a practical guide for researchers looking to incorporate LLMs into their research on text analysis. We find that, when provided with detailed label definitions and coding examples, an LLM can be as good as or even better than a human annotator while being much faster (up to hundreds of times), considerably cheaper (costing up to 60% less than human coding), and much easier to scale to large amounts of text. Overall, LLMs present a viable option for most text coding projects. Our software allows to easily use LLMs as annotators and it is publicly available at https://github.com/lorelupo/pappa.
keywords:
text analysis, text coding, large language models, GPT, fatherhood roles
1 Introduction
Political science is among the social sciences that have most eagerly embraced automatic text analysis methods grimmer2021machine (22). Indeed, in many downstream research applications, such as data collection, topic classification, and especially content analysis, political scientists have found natural language processing (NLP) techniques to hold significant potential in terms of cost-saving and scaling capacity. Various applications (text scores, topic models, etc.) have been introduced and, in many cases, adapted to the specific needs of the field slapin2008scaling (43, 21, 38). Yet, despite their promise, data scarcity (the lack of sufficient high-quality training data and the cost and time associated with creating them) and a lack of technical expertise have thus far prevented NLP methods from being widely implemented—both in political science and in the social sciences more generally.
Recently, a new generation of large language models (LLMs) has changed the cost-efficiency calculus for political scientists interested in the possibilities of NLP techniques. LLMs, including the notorious GPT series and Bard, can not only produce coherent texts almost indistinguishable from those of human writers, they can also (at least partially) automate many tasks—including coding text for analysis. LLMs need only a few examples written in plain language to learn a coding scheme, a property called “few-shot learning” brown2020language (8). This is possible because the model has a general concept of language and only needs to be specialized on the task at hand, a significant improvement over both traditional statistical text classification approaches and the most modern transfer learning ones laurer_van2023_trans (31), which require thousands of hand-coded data points. In addition, web interfaces like ChatGPT and APIs allow the use of LLMs with little programming knowledge. [^1] These models are called “closed” because their inner workings are not known. Several “open” language models exist; however, they typically require some amount of programming knowledge to use. In this paper, we focus on GPT models with a web interface. In the future, companies like HuggingFace will likely provide access to many open models via an API. These properties make LLMs attractive for political scientists who work with text but might not have the computational knowledge or resources required for other approaches. LLMs represent a shift in the potential scale, cost-efficiency, and broad accessibility of text analysis. It is likely to change the type and volume of research done, as well as the methods and the volume of data used (and hence the reliability of the results). This is the case not only for the technical fields in which LLMs were developed (computer science, AI, NLP, and machine learning) but also in political science, higher education, and various other disciplines.
In this paper, we use a concrete political science project to illustrate the workflow of text annotation with LLMs. Specifically, we code “fatherhood roles” in Swedish political discourse (see Section 3 for details). The different fatherhood roles we define are theoretical constructs with subtle differences from one another, making the coding task harder than many typical NLP coding tasks like sentiment analysis, where human annotators typically have high agreement because of the net difference between the labels. We define three different, interdependent coding tasks of varying difficulty, which we then use to code an original Swedish policy-related document dataset from 1993–2021. We annotate the data with trained human coders, GPT-3, and GPT-4.222We do not cover open-source models like LLaMa touvron2023llama (44), as they require specialized programming knowledge to use. However, our open-source software includes the possibility to use them as annotators. We also experiment with and guide the reader through some relevant choices in the LLM-coding methodology (Section 4.5): joint versus disjoint coding of multiple interdependent tasks, level of detail in the prompt description of labels, zero-shot versus few-shot learning, and the order of coding examples provided in few-shot learning. For each of these settings, we compare the performance of the language models to that of human coders along the various tasks, evaluating the quality, cost, and speed of both human coders and machines. Importantly, our evaluation is not subject to “data contamination” golchin2023time (18) because we can guarantee that the LLMs have not encountered the tasks we are evaluating during their training. Contamination arises when models are trained on popular, open-source text datasets that may include the very tasks on which they are evaluated, and potentially even the specific text examples that are present in their test set. Such exposure during training leads to models essentially memorizing correct responses, resulting in inflated performances on the “contaminated” test sets. This phenomenon affects popular datasets across various scientific fields, including political science, casting doubt on the reported performance of models’ evaluated on them. While current LLMs might have read the Swedish documents collected in our dataset during their training phase, the coding tasks are surely novel to them and cannot have been memorized.
To summarize, our main contributions are:
  1. 1.
    We find that individual coders have widely differing label outcomes, which do not always agree with others.
  2. 2.
    Evaluating two versions of GPT (3 and 4) on three political science coding tasks, using different prompting strategies, we find the best way to use the models is to provide them with comprehensive label definitions and several coding examples (i.e., few-shot learning).
  3. 3.
    In this setting, GPT-3 does not quite reach human performance, but it is comparable. GPT-4, in contrast, consistently outperforms human coders, agreeing more with each individual than the humans do with each other. Both LLMs are orders of magnitude faster than any human coder.
  4. 4.
    We find that jointly performing three coding tasks on the same texts results in comparable performance to performing them separately; the former approach is also approximately three times cheaper and faster.
2 Language Models in Political Science Research
AI, with its promise of harnessing an abundance of data, has been eagerly adopted within political science grimmer2021machine (22). Using various NLP and machine learning (ML) techniques, “computer assistants” have been used to read and accurately assess political texts in a fraction of the time it would take human research assistants to perform the same tasks. These techniques, which heavily rely on ML models benoit_2020text (7), have allowed for sophisticated analyses, from descriptive tasks such as understanding the delegation of power in the EU anastasopoulos_2020understanding (4) to more interpretative assignments, such as topic classification wang_2023_trans (46) and the surveillance of vast amounts of text to detect and flag political conflict with considerable accuracy hu_2022conflibert (27). Computer assistants have proven their significant potential for political science.
However, considerable thresholds for outsourcing downstream research tasks to computer assistants remain. Typically, ML models require large amounts of manually annotated data to train for the assignment. And once trained, they are not easily adapted to another assignment. Combined with the technical expertise required to maneuver these models, this data scarcity creates considerable barriers to implementing computer assistants, begging the question of whether switching from human to machine is worth the effort.
The data scarcity issue has triggered research dedicated to transfer learning osnabrugge_2023trans (34, 31): training the model on a dataset similar to what you want to label, thus relaxing the requirements for original annotated data.
The recent launch of Generative Pretrained Transformers (GPTs) can lower the barrier to entry even further. Developed by OpenAI, GPT language models radford2018improving (36, 37, 8, 32) can be accessed through web interfaces such as ChatGPT, OpenAI’s Playground, or programmatically through an API. In 2020, OpenAI presented GPT-3 brown2020language (8), which boasts 175 billion parameters and underwent training on a vast dataset of over 45 terabytes of text. This setup makes it one of the biggest generative language models ever developed, only surpassed by a handful of other LLMs—including its younger sibling GPT-4, the most recent and powerful model deployed by OpenAI openai2023gpt4 (32). GPTs have several practical advantages over other LLMs, particularly over a family of language models commonly used in coding tasks: BERT models.
The first is that GPT can be queried through an intuitive interface and does not require any fine-tuning (the neural network remains unchanged for whatever coding task), only proper instructions (prompting). Most BERT-style models must be loaded on your PC or server. The necessary software needs to be installed and then run on thousands of examples to fine-tune the model to the coding task at hand, a task that typically requires powerful computers with GPUs.
The second advantage is that, unlike all but a very few traditional language models, GPT can work with multiple languages. Moreover, GPT can handle a long set of instructions and examples. In contrast, BERT models need the classification task to be expressed in one sentence or a little more, making them unsuitable for classification tasks requiring accurate description. Furthermore, BERT models cannot simultaneously code a text across multiple tasks, as we show to be possible with GPT.
In economic research, (korinek_language_2023, 29, ) demonstrates 25 scenarios where LLMs like GPTs can automate tasks like ideation, analysis, and software development. Empirical studies have explored LLMs’ ability to mimic human behaviors in surveys and experiments. Notably, (argyle_out_2023, 5, ) revealed GPT-3’s capacity to replicate diverse survey responses, and (aher_using_2023, 2, ) successfully used LLMs to simulate human subjects in research, reproducing key social science findings. Extending this, (horton_large_2023, 26, ) showed GPT-3’s effectiveness in economic gaming simulations, while dillion_can_2023 (13) found GPT-3.5’s moral assessments align with human judgments. Furthermore, rosenbusch2023accurate (40) discovered GPT-3’s accuracy in generating hypotheses about social phenomena, highlighting the potential of LLMs in research automation.
These pioneering investigations merely scratched the surface of LLMs’ potential to streamline research processes. Unsurprisingly, LLMs show great potential in automating text coding assignments ornstein2022train (33). In this article, we contribute to this line of research by illustrating how language models can effectively automate and apply a novel theoretical framework to code original empirical data.
3 The Example Case Study
As a case study for the proposed methodology, we investigate how fathers are discussed in Swedish policy documents, i.e., government investigations, propositions, motions, and other legislative texts. The background of the study is the emergence of a “new father”—the idea that societal discourse is becoming less dominated by a traditional fatherhood role, and more open to other alternatives—and the debate, in academia and elsewhere, about whether and when such a shift occurred, the content of the shift, and its consequences. The study uses data from Sweden over the period 1993–2021. The substantive research question—whether fathers have been assigned a more agentic fatherhood role over time in these documents—will be elaborated on in a forthcoming article.
Stereotypical fatherhood roles in texts are more or less always analyzed using qualitative interpretive analyses graf2023post (19, 20). This means that political science has developed valuable theoretical and normative concepts and understands specific texts and cases empirically dermott2015more (12, 23, 35). However, there is a dramatic under-usage of quantitative text analyses and few possibilities for concluding whether fathers are recognized differently in different types of documents, over different periods and contexts, and by different types of message senders. For example, it is highly unclear whether societal norms around fatherhood are changed via elected representatives (so that norm change is found in motions, speeches, or the like) or via public officials and administrations (public investigations), or decided by who ends up in government (propositions, government regulations, etc.). Furthermore, the absence of quantitative measures in the field renders it unclear whether the different fatherhood roles are common or uncommon in these texts. [^3] For a recent study using experimental designs to study fatherhood roles, see (scheifele2021representations, 42, ) Thus, this case study represents a political science research area where large-scale automatic text coding is needed.
3.1 Data
We used publicly available data from the Swedish parliament (Riksdagen). [^4] https://www.riksdagen.se/sv/dokument-och-lagar/riksdagens-oppna-data/ We collected all public documents from 1993 to 2021, converted them to raw text format, and lowercased them. We selected only sentences referencing fathers (i.e., containing the Swedish words pappa, pappor, fäder, fader, far). We added further checks to ensure the term actually referred to a father: for example, the word “far” can also mean “go” or “travel” if used as a verb. We used a part-of-speech tagger to filter out those instances. The final data set contains 1,910 sentences. Data collection and filtering details are reported in B.
4 Methodology
In this section, we present the main steps researchers should follow in an LLM coding task after collecting their data:
  1. 1.
    Coding paradigm: deciding what the labels are.
  2. 2.
    Validation set: hand-code a few hundred examples to evaluate the LLM’s performance.
  3. 3.
    Language model: choose one, or compare multiple models before selecting the most suitable.
  4. 4.
    Prompt: construct the textual input to the model, including a concrete instruction, a description of the labels, and a few hand-coded examples (which should not be part of the validation set).
  5. 5.
    Prompt engineering: fine-tune the prompt by trying out different options. For each option, let the LLM code the validation set and evaluate its label predictions against human labels.
  6. 6.
    Coding the entire corpus by querying the LLM with the best-performing prompt obtained from the previous step.
4.1 Coding paradigm
After an initial qualitative analysis of our dataset, we classified each of the Swedish texts on fatherhood along three tasks, each represented by a varying number of suitable labels:
  1. 1.
    Type of paternal involvement in the family. Possible labels for this task are “passive,” “active negative,” “active positive caring,” “active positive daring,” “active positive other,” and “not applicable.”
  2. 2.
    Explicitness of the description, denoting the linguistic means by which the father’s role is expressed in the sentence. Possible labels: “implicit” and “explicit.”
  3. 3.
    Normativeness of the description, denoting whether the sentence expresses a factum or a desired outcome. Possible labels: “descriptive” and “ideal.”
The three labeling tasks target different aspects of the texts (content vs. form), are interconnected, and are hierarchical. If the first task is “not applicable”, the other two tasks are irrelevant and thus not coded. B presents a detailed description of each label.
4.2 Validation set
Three authors independently coded the same random subset of 350 sentences, applying the coding paradigm described above. This validation set allows us to establish a baseline of human performance. It gives us a direct comparison set for the coding choices by the LLMs. This step is optional but highly recommended, as it is needed to measure the performance of the LLMs and to carry out the prompt engineering step to maximize performance.
4.3 Language Model
For most of our experiments, we used the GPT-3 model from OpenAI via their API. [^5] https://platform.openai.com/docs/api-reference/introduction. Specifically, we used the text-davinci-003 language model to classify texts into our predetermined labels. [^6] https://platform.openai.com/docs/models Due to the model’s language pretraining, it is sophisticated enough to learn the labeling task from just a prompt, the components of which will be discussed in the following section. Following the best prompting strategy identified in the prompt engineering step, we also evaluated the most powerful (and most expensive) OpenAI model available to date: GPT-4. However, even for that model, the overall cost is moderate (see Section 5.3).
4.4 Prompt
For GPT, we used the exact same instructions and label definitions as would be used in the codebook for human coders. However, how they are provided to the model differs along some dimensions.
Instructions
Language models need clear instructions that explicitly describe the goal of the task. Instructions can be formulated in any of the languages understood by the model. However, performance might vary according to the languages used. We used English instructions, even though the text coded by the model is in Swedish. In fact, in preliminary experiments we found that Swedish instructions alongside Swedish texts resulted in inferior performances compared to English instructions. The most likely reason for this behavior is that GPT-3 is mainly trained on English data brown2020language (8), and, therefore, English is the language it understands best. This finding suggests that GPT should be a stronger coder of English texts than non-English ones, as corroborated in other recent studies ahuja2023mega (3, 14).
The instruction we used for coding all tasks reads as follows:
Label the Swedish text according to how it describes the role of the father in the family. Possible labels are:
Labels
We next provided the codebook to the language model prompt. Besides providing the model with the list of labels defined in the previous coding paradigm step (Section 4.4), the model can also benefit from a description of these labels. Depending on the labeling task, the labels might be already self-explanatory (such as for task 2) or unclear (such as for task 1). As we will see in the experimental section, in the latter case an exhaustive description, including ample formulation of the details and hints to identify relevant texts, usually works better than a synthetic description, which in turn is better than no description at all. Our codebook, including labels and their definitions, can be found in B.1. In the experiments, we simply copy-pasted this codebook into the prompt after the instruction.
Example sentences
Following the codebook, we provided GPT-3 with 15 example sentences from our corpus that we manually labeled with the three desired labels. GPT-3 used these examples to learn how sentences map to labels and apply this mapping to the texts given to it. It is important to include examples for all of the labels, and that these examples are not part of the validation set. We randomly sampled these examples from our corpus, excluding the validation set. The examples used in our experiments are reported in C.1.
The procedure of providing a handful of examples in the prompt is called few-shot learning. The selection of these examples can significantly impact the coding performance of the language model. For this reason, we recommend measuring the model’s performance on the validation set when provided with different sets of examples. Note that each example is counted against our word budget, with more words meaning higher costs (see Section 5.3).
Alternatively, we can perform zero-shot learning by leaving out all examples and providing only the instructions and label descriptions. This is much faster, less prone to incorporate human biases, and potentially cheaper, but known to be potentially less accurate.
4.5 Prompt engineering
To optimize GPT’s coding performance, we evaluated results from different prompting combinations, varying the following features of the codebook and the example sentences:
Tasks
Indicates whether the model is prompted to code a single task at a time, or the three tasks simultaneously, for each text that is provided in input. In the former case, the model outputs a single label for each text, while in the latter it generates three, one for each task. For example, in Table 1, we prompted the model to code task 1 alone, except in the last two rows, were we ask to code all tasks. The idea behind the second option is to use shared information in the hierarchical coding scheme, and to make the coding process more efficient, as the model does not have to be run three times on the same instance to provide a label for each task.
Label descriptions
Indicates different degrees of explanation of the labels provided to the model, from a long and comprehensive explanation to a concise one to none. We experimented with all three variants to code task 1 (type of paternal involvement), where labels have complex semantics. In contrast, we coded the second (explicitness) and third (normativeness) tasks with self-explanatory labels. Further explanation is in these cases redundant and hence not provided to the model.
Number of example sentences
We experimented with both few-shot and zero-shot learning. In the former case, we provided 15 labeled Swedish text examples to the model after the instructions and the definition of the labels. In the zero-shot case, we provided none.
Examples order
In D, we investigate how different orders of the example sentences provided in the prompt affect coding quality. Although the effect is not significant, we suggest strategies to avoid “unfavorable” orders.
5 Evaluation
We evaluate the automatic coding along three variables: quality, speed, and cost.
5.1 Quality
One advantage of letting GPT-3 label text instead of humans is the higher chance of getting consistent coding over time. Humans are known to be inconsistent (they become hungry, get a text message, need to go to the bathroom) and often disagree with their previous assessments when queried again abercrombie-etal-2023-temporal (1). They also frequently and systematically disagree with other human coders, even if they have received the same training and instructions. Coding using GPT-3 should be more consistent, as long as temperature is kept to 0 (this parameter controls how “creative” GPT’s answers are).
To investigate the reliability of humans compared to GPT, we first compare the performance of three human coders (all authors of the paper) to each other on the validation set. We evaluate their coding along three measures: chance-corrected agreement (Cohen’s $\kappa$), raw agreement (Raw), and predictive performance (F1). Cohen’s $kappa$, a measure of inter-annotator agreement, is usually between 40 and 60 landis1977measurement (30). Less-than-perfect results for the human coders do not necessarily signal bad training or bad coders but usually indicate how difficult a task is and the absence of a ground truth. To measure the raw agreement of an annotator, say “Human 1,” with the others, we use the other human annotators’ set of labels as the gold standard and measured how often Human 1’s label predictions matched them on average. The predictive performance is calculated similarly, but accounting for class imbalances (i.e., gold label “passive” being more present in the test set than the label “active_negative”). The results of this evaluation can be found in the first block of rows of Table 1, 2, and 3, where the first, second, and third labeling tasks are evaluated, respectively.
Besides providing the performance of each human annotator, we also measure their average performance (“Humans AVG”).
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 49.30 61.29 57.22
Human 2 n.a. n.a. n.a. 51.02 63.86 57.60
Human 3 n.a. n.a. n.a. 51.65 64.29 57.54
Humans AVG n.a. n.a. n.a. 50.66 63.15 57.45
GPT-3 1 none 0 13.41 27.71 22.73
GPT-3 1 short 0 20.89 36.38 34.27
GPT-3 1 none 15 36.53 51.05 44.22
GPT-3 1 short 15 44.50 58.48 49.40
GPT-3 1 long 15 49.36 62.57 54.92
GPT-3 1,2,3 long 15 48.34 62.48 54.26
GPT-4 1,2,3 long 15 54.05 66.38 61.70
Table 1: Average coding agreement with (other) human coders on task 1: type of paternal involvement (6 labels). We evaluate both human and GPT performance with different combinations of the prompt in terms of: annotation Tasks performed simultaneously, length of the Label descriptions, and N. of Examples provided to the LLM.
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 26.74 54.86 50.13
Human 2 n.a. n.a. n.a. 26.55 58.29 48.61
Human 3 n.a. n.a. n.a. 29.75 59.14 52.33
Humans AVG n.a. n.a. n.a. 27.68 57.43 50.36
GPT-3 2 none 0 14.66 42.76 34.72
GPT-3 2 none 15 22.91 53.14 48.54
GPT-3 1,2,3 none 15 25.91 59.14 50.07
GPT-4 1,2,3 none 15 31.97 63.62 53.49
Table 2: Average coding agreement with (other) human coders on task 2: explicitness of the description (2 labels).
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 41.64 66.57 61.83
Human 2 n.a. n.a. n.a. 37.00 67.14 57.94
Human 3 n.a. n.a. n.a. 44.17 71.14 63.20
Humans AVG n.a. n.a. n.a. 40.94 68.28 60.99
GPT-3 3 none 0 7.35 40.29 29.94
GPT-3 3 short 0 7.96 36.86 32.00
GPT-3 3 none 15 39.75 68.76 60.97
GPT-3 3 short 15 44.08 70.86 64.25
GPT-3 1,2,3 short 15 41.85 70.48 62.67
GPT-4 1,2,3 short 15 40.55 71.81 61.43
Table 3: Average coding agreement with (other) human coders on task 3: normativeness of the description (2 labels).
The task is sufficiently difficult even for trained, motivated, and expert human coders. The average Cohen’s $\kappa$ indicates “substantial” agreement, though on the lower end of that designation. The takeaway is that annotating fatherhood roles is tricky.
Using zero-shot learning (i.e., zero examples) results in bad performance across the board. Longer, comprehensive label definitions help whenever they are relevant, i.e., for tasks one and three. However, they cannot match the improvement we see with 15-shot learning, even without an explanation of the labels. The combination of few-shot learning and long, comprehensive label definitions results in the best performance, comparable to that of human coders and sometimes even outperforming them, as was the case for our third task.
Interestingly, GPT-3 performs similarly when labeling the three tasks jointly and separately, in different runs. Performance on the second task is slightly better with joint labeling, while the contrary is true on the third task. This is good news because jointly labeling across the three tasks is faster and cheaper: we only need to prompt the model to label each text once instead of three times.
GPT-4
When prompting GPT-4 with the best prompting strategy found in our experiments, it consistently outperforms GPT-3 and all individual human annotators. The fact that GPT-4 agrees more with individual human annotators than they do among themselves might appear counter-intuitive. However, we should remember that the models are coding the texts based on 15 examples whose labels all the human annotators agreed upon. Therefore, the models’ coding is heavily influenced by coding choices that are shared by annotators and reflect this agreement.
5.2 Speed
To evaluate the speed and cost of automatic coding, we let GPT-3777We choose GPT-3 over GPT-4 for this evaluation because it has higher rate limits. At the time of writing, GPT-4 was a brand-new technology, and its production was not yet mature. Therefore, GPT-3 was more representative of the speed and costs that the researchers will incur in the near future, when GPT-4 is expected to expand rate limits and lower costs. annotate our entire dataset of 1,910 sentences with the best prompting strategy resulting from the previous evaluation step, i.e., using 15-shot learning, providing a long, comprehensive description of the labels, and prompting the model to annotate across the three tasks jointly.
The completion time for GPT-3 was about 16 minutes. [^8] This time can vary slightly depending on how busy OpenAI’s API is. Meanwhile, based on other projects we have run, and ignoring diminishing returns due to fatigue, a trained and motivated human coder can label about 100 sentences of this type per hour. A single coder would thus take about 19 hours, or at least three work days, including the necessary breaks and excluding separate training sessions, adjudication of difficult cases, etc., to label our dataset. If coding decisions became more complex, this time estimate would increase.
Thus not only does GPT-3, unlike other automatic coding approaches used in the literature, require no training time, it is also approximately 270 times faster than a human coder.
5.3 Cost
The total price of a labeling request to OpenAI’s API is computed based on the number of words in the prompt and response. [^9] https://openai.com/pricing When coding a corpus like ours, the instruction, label definitions, and labeled examples present in the prompt are fixed across requests, while the sentence to be labeled changes at each API request. Since the sentences to be coded are of different lengths, the price for coding each sentence varies. The price of the response depends on the labels chosen by GPT-3, but it is negligible since each consists of only a few words. In total, we spent \$93 coding the 1,910 sentences in our data, an average price per instance of around 5 cents.
A realistic estimate for a human coder in the U.S. is a student assistant salary of \$12.35 per hour. [^10] We use US standards for the cost of human coders since the pricing for GPT-3 is in US dollars Assuming the coder takes 19 hours to label the data, the total cost would be \$234.65. That is more than 2.5 times the price of the LLM. If we wanted multiple coders, as we simulated here, the overall price would go up, in our case, to a total of \$703.95.
In political science, we are often interested in country-specific text analysis, requiring native-speaking coders with appropriate knowledge of the field. In our case study with Swedish data, for example, this would incur a salary of about 170 SEK per hour, amounting to about \$881.79 for 19 hours of work for three coders, excluding recruitement costs.
As the LLM economy develops, APIs providing language modeling services are expected to become faster, cheaper, and more powerful, making LLMs an even more economic choice.
6 Discussion
Our results show that using an LLM for text coding is considerably faster and cheaper than recruiting, training, and evaluating human coders. It also achieves matching predictive performance and, in several cases, even performs better than individual human coders. Moreover, like human annotators, the LLM can annotate texts along multiple coding tasks without losing quality, and even improving it. We expect the outcome to be even better for English text coding and with languages that are more abundant online (such as Spanish or German), as they are more present in the language model training data openai2023gpt4 (32).
Of course, we can retrain human coders to improve their agreement with each other and their label accuracy, possibly outperforming LLM coding. However, that process requires even more time and incurs more costs.
But do we want to replace human coders with LLMs completely? That is a different question. LLMs are known to hallucinate (i.e., fabricate facts) and suffer from alignment issues (i.e., their moral underpinnings). Indeed, recent research has shown that different LLMs show different political leanings feng-etal-2023-pretraining (15). Hallucinations can be addressed by mapping non-matching labels to a default (for example, “not applicable” or the majority label), and repeated runs can reduce their risk. Moreover, the concern about hallucinations is larger when generating blocks of text. As we only request individual labels and give very explicit instructions, this problem is unlikely. Moral alignment is a trickier problem, although it can be solved for some tasks with adequate prompting argyle_out_2023 (5). These facts together make LLMs a sensible choice for sensitive data and for morally ambiguous or politically charged tasks. In these cases, constructing a validation set is suggested even more strongly to verify the alignment of LLM answers with human coding before deciding whether to annotate more data.
Labeling by LLM also offers another possibility, namely freeing up (human) time and resources to check the reliability of the coding. Too often, we assume that trained human coders are reliable and consistent, despite a growing body of evidence that this is not the case abercrombie-etal-2023-temporal (1). However, by outsourcing the first round of labeling to an LLM, we can use the time otherwise devoted to coding to check for inconsistencies, interrogate differences, and perform additional analyses.
7 Conclusion and Outlook
“If you want something done right, do it yourself.” For the longest time, this statement has highlighted the dilemma between efficient use of resources on the one hand and quality on the other. For a while now, the prospect of AI research assistants has promised to resolve this dilemma. But despite high quality, the technical expertise and vast training data requirements to operate these models have constituted severe thresholds for delegating to computers—delivering on promises of quality but eroding the prospects of resource efficiency. But as LLMs have been deployed, exhibiting outstanding results in cognitive tasks from economic research, writing, and data analysis (korinek_language_2023, 29, ) to acing the U.S. Bar exam and the Law School Admission Test (LSAT), they begged the question: If we want something done not only right but quickly and cheaply, is it finally time we let the computers do it?
In this paper, we apply LLMs to a typical downstream research task: data annotation. We present a case study on using LLMs (here, GPT-3 and GPT-4) to automate text coding, provide a general framework that can be applied to any text classification task, and discuss essential parameters and their effects on the outcome. Finally, we evaluate the quality of human-generated annotation compared to computer-generated annotation.
Using LLMs for text coding has several advantages over the traditional approach via either trained coders or supervised classifiers. Using LLMs is considerably faster (in our experiments, more than 270 times) and cheaper (at least 4 times) than relying on human coders. Regarding quality, GPT-3 (the cheaper alternative) performs on par with the individual human coders, while GPT-4 outperforms them. Though high quality and similar time gains can be achieved using supervised BERT-based classification models laurer_van2023_trans (31), LLMs require no specialized programming, can be implemented relatively simply via an API, do not require thousands of pre-coded examples to train the model, and can work with multiple different languages. By matching the results of individual human coders on a Swedish dataset with a seed set of just 15 coded examples, we show that LLMs are valuable for political scientists interested in text analysis in many languages.
An essential issue when working with LLMs is their prompting. This paper investigates the thoroughness of label definitions, the number of shots and the number of tasks included in the prompt. We find that all three aspects affect the performance of the LLMs. Few-shot learning has the most significant effect, but label definitions also improve performance. We also show that while annotating several tasks simultaneously has no clear reliability improvements, neither is the reliability lowered. This finding suggests that multiple interdependent annotation assignments can be performed at the same time and price by the LLMs without a reduction in quality but with an increase in cost-efficiency.
While our case study has limited scope, it adds to the growing body of evidence around LLMs as useful annotators. Our results suggest that if we want something done right, we might as well have an LLM do it.
LLMs will likely have an outsized impact on political science methodology, not just in text analysis but also in text generation, such as survey questions, use of agent-based design, and manuscript outlines. The technology is here to stay, and we hope this paper helps practitioners to make sense of LLMs and adopt them responsibly and effectively for their needs.
Funding.
This work was supported by Riksbankens Jubileumsfond, project number RRD10-1418:1; Knut and Alice Wallenberg Foundation, project number KAW 2017.0245.
Acknowledgments.
We would like to thank our colleagues at the Bocconi University’s MilaNLP Lab for their constructive feedback on several versions of the manuscript.
Data Availability Statement.
Replication code and cleaned data are available on GitHub: https://github.com/lorelupo/pappa.
Competing interests:
The authors declare none.
Appendices
Appendix A The evolution of language models
OpenAI’s GPT is one of the most advanced language models available to the public. Like all its current competitors, it builds probabilistic and neural language models. It is, therefore, useful to understand both families of models and how they work to get the most out of them. This section represents a relatively technical introductory overview of these topics.
A.1 Probabilistic Language Models
Language models have been around for several decades. In their simplest form, they are a function that assigns a probability to any sentence $S$. That probability, $P(S)$ can be understood as the likelihood of encountering the sentence “in the wild.” Language models have been used to assess the output of various applications, for example, machine translation, and to re-rank the output based on their likelihood.
The most common way of estimating a probability is via maximum likelihood estimation (MLE)—or counting and dividing on a training corpus. The resulting probabilities can then be used (in theory) on any future input. Here, we would count how often we have observed the sentence in a training corpus and divide it by the number of all sentences in that corpus. One obvious obstacle to this approach is that there are countably infinite possible sentences (any sentence could be extended by adding another sentence via “and”). The normalization factor is, therefore, infinity. Since dividing any number by infinity gives the answer 0, this approach does not work. Thus, language models have to be able to assign a probability to unseen sentences.
For the longest time, language models were based on the maximum likelihood estimation of $n$-grams. That is, they computed the probability of a sentence as the joint probability of all its words: $P(S)=P(w_{1},w_{2},...,w_{n})$. Again, the joint probability is hard to estimate directly. However, it can be broken down further into a sequence of conditional probabilities of word combinations (the $n$-grams).
So, instead of computing the probability of “I like platypuses,”, a 2-gram language model computes the probability of I at the start of a sentence, the probability of like given that the previous word was I, and the probability of platypuses given that the previous word was like. Or
$P(I,like,platypusses)=P(I|\texttt{START})\cdot P(like|I)\cdot P(platypusses|like)$
The first two can be computed by taking a large enough corpus and counting how often we see I at the beginning of a sentence, and dividing by the count of all words starting a sentence ($P(I|\texttt{START})$). We can then count the number of times we observe I like and divide by the count of any I followed by any word. The last word, platypuses, is unlikely to occur in a corpus. Probabilistic language models use smoothing to ensure that a non-zero probability is assigned to every word in the vocabulary, even the rarest. The simplest form is Laplace smoothing, which adds a constant to the count of any word. To still produce a valid probability, we must recompute the normalization factor. There are various ways to do this.
Probabilistic language models of this type can be implemented in a few lines of code and assign reasonable probability estimates. Increasing the size of the $n$-grams produces better estimates. Still, it requires larger and larger corpora and more sophisticated smoothing chen1999empirical (9).
Instead of scoring sentence probabilities, these models could also be used to generate new sentences by sampling from the learned conditional probability distributions. I.e., for each word, we select one from the most likely ones we have observed in that position. The resulting sentences are locally coherent (i.e., any three words in a row would make sense) but often lack cohesion as they get longer and have no coherent meaning across sentence boundaries.
A.2 Neural Language Models
Probabilistic language models were limited by their reliance on $n$-grams and the Markov horizon to compute the necessary statistics. Larger $n$-gram windows produced better models but required larger corpora and more storage and improved only incrementally. In the 2010s, a resurgence in the use of neural networks allowed for a dramatic leap forward in language models. Neural networks had been around since the 1950s rosenblatt1958perceptron (39), but interest in them died off in mainstream AI as they were too inconsistent in their results. A variety of factors led to their resurgence.
The main factor is the availability of unprecedented amounts of data, driven by the popularization of the world wide web and, more recently, mobile applications. Neural networks have large numbers of parameters, and to set these parameters correctly, they need to be trained on large amounts of data. When IBM’s Watson beat the leading Jeopardy! players by a wide margin in 2011, it used a data dump that comprised much of the collected knowledge available on the Internet from the year before. But the underlying technique, having independent algorithms come up with solutions and then weigh and combine them, was 20 years old. The technique had always been viable, but not without the vast amount of data that had only recently become available. Indeed, a comparison of the most common AI problems shows that the algorithm that “solved” each of them was typically invented 18 years before the problem was solved. [^11] https://www.kdnuggets.com/2016/05/datasets-over-algorithms.html Meanwhile, the data sets that led to the breakthroughs were, on average, published just three years before.
Concurrently, researchers have made progress in understanding how these parameters should be computed and controlled, reducing the variability in neural network results. The development of better initialization regimes glorot2010understanding (16), better non-linear activation functions glorot2011deep (17), and an improved understanding of exploding or vanishing gradients hochreiter1991untersuchungen (24) has made it possible to train neural networks more efficiently and reliably.
The final factor that led to the resurgence of neural networks was computing power. Moore’s law schaller1997moore (41) predicted doubling the number of transistors on a microchip every two years. While that relation has slowed, the computing power available on a simple smartphone today is magnitudes larger than anything available at the dawn of AI and even as recently as 20 years ago. Computations that would have taken years to finish on a specialized supercomputer can now be run on a standard laptop. Specialized hardware like graphical processing units (GPUs) has become even more specialized and is the architectural backbone of today’s neural network models.
Neural networks, despite their names, are essentially universal function approximators. They can learn to replicate any function, provided they have enough data. This includes probability functions. Neural networks can directly compute the probability distribution over all words for any sequence without breaking it down into a sequence of conditional probabilities or limiting the horizon. This has made it possible to produce much longer coherent sentences with them.
A.3 The Transformer Architecture and GPT
Until recently, neural networks struggled with long-range dependencies, such as discourse phenomena in text sequences. In the sentence “This is not, by any stretch of the imagination, a funny movie,” the keywords not and funny are far apart, yet need to be connected to make sense of the whole. In 2017, (vaswani2017attention, 45, ) introduced a groundbreaking neural architecture called Transformer that surpassed previous architectures used for sequence modeling, such as recurrent hochreiter1997long (25, 10) and convolutional kim-2014-convolutional (28, 11) neural networks. At the heart of the Transformer architecture is the attention mechanism bahdanau2014neural (6), a neural module allowing the elements of a sequence to be linked directly to one another, even when distant. Transformer-based language models are today state-of-the-art and have gained public attention thanks to the remarkable linguistic prowess displayed by GPT models.
Until recently, neural networks excelled in specific tasks, such as sentiment analysis, but they couldn’t bridge the gap with human-like versatility. This gap is narrowing with the latest LLMs like GPT, which demonstrate a broader skill set, including software development, translation, mathematical derivations, and rhyming.
Appendix B The corpus and codes
Policy documents here refer to official texts written in everyday work by or for the Swedish Riksdag and Government. Our corpus includes various policy documents, for example all the motions, propositions, written questions, and interpellations published during the period 1993–2021, further described below. The policy documents have in common that they all somehow preceded parliamentary decisions and include arguments for why a law should or should not pass. For instance, the Swedish lawmaking process is preceded by so-called public investigations via the SOU system (Statens offentliga utredningar). Going back in time, all laws have a corresponding “SOU: year: number” where skilled investigators provide the lawmaker with arguments for laws. It is likely that motions pertaining to the agentic role of the father that actually have had an impact on lawmaking (and therefore possibly on the behavior of men) would be found here. Furthermore, it is likely that government propositions and motions introduced by political parties and individual MPs will contain references to the role ascribed to fathers by lawmakers, i.e. parties who currently constitute the government as well as those in the opposition in Parliament.
We downloaded the policy documents from the website https://data.riksdagen.se/data/dokument/, which makes Swedish Riksdag and Government documents available as a publicly accessible dataset.
The Swedish names for the documents we include are: betänkanden, departementsserien, interpellationer, motioner, propositioner, protokoll, skriftliga frågor, statens offentliga utredningar (SOU), utredningar, utskottsdokument, yttranden, övrigt. We downloaded 16,390 documents under these headlines dating back to 1961. Most of them were motions (7,361), protokoll (3,263), betänkanden (1,549), propositions (1,244), and SOUs (1,116). Of these, we used only documents from 1993 and onward in this paper, as we cannot be sure that the documents available online for earlier periods represent complete sets of all documents of each type in those periods.
The documents were processed in the following way: we downloaded all available documents, linked the content to the file ID and type, and extracted the data from PDFs where necessary. We lightly preprocessed the data (lowercasing and replacing numbers), then split it automatically into sentences, separating words and punctuation marks from each other. Thereafter, we identified target words for fathers, mothers and parents.
There were 277,168 sentences that included any of these keywords. We singled out the instances where these words were used as nouns, and not as verbs or other types of words. This means that our analyses take into account, for example, that “far” (father when used as a noun) is also a verb meaning “go” or “travel.” This left us with 151,489 sentences that include the keywords as nouns. The overwhelming proportion of these sentences talk about parents and do not single out fathers or mothers as separate actors: 1,901 sentences mention fathers and 2,821 mention mothers.
The aim of our annotation is to detect the role that the father plays in the life of the family. We divided the annotation into the labels described in the following codebook.
B.1 Codebook
PASSIVE
This is a category that captures what is sometimes called the classical role given to and taken by fathers, which in essence is a role that does not include engaging with the child in person. Words used in theory to describe this role might be “custodial parenthood,” “provider,” “breadwinner,” or “process parenthood.” This is a person who is important in the life of the child, but is not involved in their hands-on care and upbringing. Society includes norms and sometimes creates clarifying policies around this father, for his sake, as well as for the sake of the child and the mother. Envisioned indicators might be words denoting the father being the moneymaker or provider, or someone who provides genes or sperm or overall protection or security.
ACTIVE NEGATIVE
This is a category that captures the negative roles of fathers, such as the oppressive father, the harsh and violent father, as well as the neglectful father. Society needs to create policies around this father in order to restrict him. Envisioned indicators are notions such as dangerous, punish (not educate or discipline in a non-aggressive way), aggressive, violent, oppressive, or not paying child support.
We also have three categories that capture the positive roles of fathers. Society needs to create policies around this father in order to encourage him. These categories are:
ACTIVE POSITIVE CARING
This category focuses on what is stereotypically seen as the female parental role. Envisioned indicators would be words such as caring, warm, nurture, understanding, empathy, listening, comfort, or confirming
ACTIVE POSITIVE DARING
This category focuses on what is stereotypically seen as the male parental role. Envisioned indicators would be things like strengthening activity, risk-taking, daring, test your limits, sport, outgoing, education (including reading to the child, which is one of the recurring activities in the data).
ACTIVE POSITIVE OTHER
There are also other ways of being active as a father, and this category captures that. Most typically, this category indicates that the father is competent and reliable, and should be trusted with the child, but it does not single out why. Envisioned indicators are general references to competence, capability, responsibility, rights of and trust in fathers to handle information, take parental leave, get parental allowance, be present in the life of the child, or be an unspecified role model (“boys/girls need male role models”). Note: If there are norms that fit the Caring or Daring categories and these are present in combination with “other” norms, these sentences will be annotated as Caring or Daring. In that way we prioritize caring and daring above the other category.
NOT APPLICABLE
Sometimes sentences include the word father but do not tell us anything about fathers, and it is not possible to detect a norm. We asked the computer to denote those as “not applicable.”
Additional labels
The sentences are also labeled either as describing a state (DESCRIPTIVE) or an ideal (IDEAL). This is because we found that fatherhood norms can be either descriptive or prescriptive in the sentences (“we see that fathers are active” or “we wish for more active fathers”). Furthermore, fatherhood roles can be either IMPLICIT or EXPLICIT in the sentences.
Appendix C Prompts
In this appendix, we provide the instructions and definitions of the labels included in our GPT3 prompts for each labeling dimension. Then, we provide the 15 examples used in few-shot learning.
Dimension 1: type of paternal involvement
Instruction:
Label the Swedish text according to how it describes the role of the father in the family. Possible labels are:
Long labels definition: exaclty the same reported in the previous appendix (B).
Short labels definition:
  • passive: fathers who are not actively involved in hands-on care and upbringing of the child;
  • active_negative: fathers exhibiting harmful behaviours like aggression, violence, or neglect;
  • active_positive_caring: fathers providing care, warmth, empathy, and support;
  • active_positive_challenging: fathers encouraging risk-taking, growth, and educational activities;
  • active_positive_other: fathers displaying competence, responsibility, trustworthiness, etc., without specifying a specific role;
  • not_applicable: not applicable.
Dimension 2: explicitness of the description
Instruction:
Label the Swedish text according to how it describes the role of the father in the family. Possible labels are:
  • implicit
  • explicit
  • not_applicable
In this case, we did not experiment with label definitions since labels are self explanatory.
Dimension 3: normativeness of the description
Instruction:
Label the Swedish text according to how it describes the role of the father in the family. Possible labels are:
Short labels definition:
  • descriptive: if the sentence describes a state, like “we see that fathers are active”
  • ideal: if the sentence is prescriptive, like “we wish for more active fathers”
  • not_applicable: not applicable.
In this case, we did not experiment with longer label definitions since they would be redundant.
C.1 Examples
The examples used in few-shot learning were the following:
Text: i båda fallen är modern genetisk mor till barnet . Label: not_applicable
Text: i de fall socialnämnden inte tar sitt ansvar eller inte lyckas , måste det ändå finnas rätt för nära anhöriga , som mor- och farföräldrar , att hos tingsrätten föra talan om umgänge . Label: not_applicable
Text: det finns en överrepresentation bland fäder med låg utbildning , låg inkomst och pappor med utländsk bakgrund . Label: passive, explicit, descriptive
Text: visserligen är det obalans i uttaget av föräldraförsäkringen , men samtidigt är det viktigt att slå fast att fler pappor i dag tar ut föräldraledighet än före införandet av mamma- och pappamånaden . Label: active_positive_other, explicit, descriptive
Text: om barnets föräldrar varken är gifta eller registrerade samman- boende vid barnets födelse är mamman ensam vårdnadshavare . Label: passive, implicit, descriptive
Text: bland dem fanns män , kvinnor och barn , unga och gamla , fäder och mödrar , söner och döttrar , makar och syskon . Label: not_applicable
Text: lagutskottet har tidigare anfört att man anser att den nuvarande ordningen ger betryggande garantier för att det verkligen är den biologiske fadern och inte någon annan som fastställs som far till ett barn ( $0$/00 : lu0 ) . Label: passive, explicit, descriptive
Text: i isf : s rapport dubbeldagar – pappors väg in i föräldrapenningen dras slutsatsen att pappors uttag under barnets första levnadsår har ökat och att pappor som tidigare valt att ej ta ut föräldrapenning nu börjat ta ut dagar . Label: active_positive_other, explicit, descriptive
Text: som förutsättning för alt barnavårdsnämnden skall vara skyldig att utreda om annan än den äkta mannen kan vara fader uppställer förslaget - utom att bamet skall ha viss anknytning lill sverige - dels att utredningen skall ha begärts av vårdnadshavare för barnet , d , v , s , normalt modern , eller av den äkta mannen , dels att det skall finnas lämpligt . Label: passive, explicit, descriptive
Text: genom projektet “ läs för mig pappa ” , knutet till arbetsplatsbibliotek , uppmuntras fäder att både läsa själva och att läsa för sina barn . Label: active_positive_challenging, explicit, ideal
Text: någon skyldighet att godkänna en utländsk dom om moderskap för barnets sociala eller genetiska mor , eller en utländsk dom om faderskap för en genetisk fars manliga partner , kan däremot inte utläsas av domstolens hittill- svarande praxis . Label: passive, explicit, descriptive
Text: här uppmanas pappor också att ta ett ansvar som läsande fäder . Label: active_positive_challenging, explicit, ideal
Text: av barnens svar kan man utläsa att mammor generellt har mer tid för sina barn oavsett i vilken utsträckning de bor med mamma , medan för barnen kan det ge mer tid med pappa när föräldrarna separerar än när de bor ihop under förutsättning att de minst bor halva tiden med pappan . Label: active_positive_other, implicit, ideal
Text: det är fullt möjligt att kommunicera kring varför för- äldraskapet så lätt upplevs som något som angår mamman mer än pappan , varför pappor kan känna sig som tillräckligt goda fäder trots ett mycket lågt uttag av föräldraledighet , varför mammor ofta tror att de är bättre på att läsa av barnet än vad pappor är , varför pappor uppfattas som hjältar när de gör samma sak som mammor gör av bara farten och tusen andra frågor som har att göra med för- äldraskapet som bekönad position . Label: active_positive_caring, implicit, ideal
Text: då utmålas oftast pappan som hotfull , våldsam och som en dålig förälder . Label: active_negative, explicit, descriptive
Text: det vanliga är sedan att våldtäktsmannen – om det inte var fadern eller en nära släkting som förövat dådet – betalar flickans föräldrar för att de inte skall anmäla fallet till polisen . Label: active_negative, implicit, descriptive
Text: för att underlätta för framför allt fäder att öka sitt uttag av föräldrapen­ning föreslås att ersättningen utges med sjukpenningens belopp oberoende av den andre förälderns sjukpenning . Label: active_positive_other, explicit, ideal
Text: ha föräldrar gemensamt barn i sin värd , utgår föräldrapenning över garanlinivån lill fadern endast under förutsättning att även modern är eller enligt vad förut sagts bort vara försäkrad för sjukpenning som överstiger garantinivån . Label: passive, explicit, descriptive
Appendix D Order of example sentences
Not only does the number of examples that we provide in the prompt affect the output of a language model, it also affects the order of the examples. The last examples in our 15-shot sample have a more significant effect on the model predictions than earlier sentences. This known the “recency effect” zhao2021calibrate (47) and is based on how the language model was trained. When predicting a missing word, the immediate context before the predicted word is more important than sentences that occurred several paragraphs before. In the following, we study the effect of this recency bias on labelling performance and a strategy to deal with it.
After constructing the prompt, we let GPT-3 label the validation set of 350 sentences by feeding it each example with the prompt via the API and collecting the answer label. We used this setup to evaluate the performance of GPT-3 against human coders (see Section 5). We evaluated different variations of the prompts to optimize for the best performances. Once we were happy with GPT-3’s performance, we let the LLM code the entire corpus, in our case 1,910 sentences.
Examples order Kappa Raw F1
0 49.36 62.57 54.92
1 48.27 61.90 53.51
2 46.58 60.67 53.72
Avg 47.97 61.71 54.05
Std 1.39 0.96 0.76
Majority 48.40 61.90 54.31
Table 4: GPT-3’s performance on the first labeling task for three different random orders of the 15 labeled sentences provided in the prompt as examples.
Table 4 shows the results of three different GPT-3 runs over the validation set, where the prompt contains a different random order of the 15 examples provided in each run. The task under evaluation, in this case, is the first and most challenging (“type of paternal involvement”). We prompted the model with only the label list for this task and long, comprehensive descriptions. As expected from previous findings on recency bias, the performance changes with the order of the examples, indicating that recency bias is at play here.
Since we know which order performs best from trying different orders on the validation set, we can use this order when coding the entire corpus. In our case, order 0 exhibited the highest reliability vis-a-vis human coding. [^12] Order 0 is also the one we used on the validation set in all the other experiments, as presented in the previous tables.
Suppose we do not have a validation set available. In that case, we can mitigate the risk of recency bias by adopting a majority voting strategy. In this case, we let the LLM code the whole corpus multiple times, with different random orders of the example sentences in the prompt, and then select for each instance the label that was output by the model most of the time, breaking ties randomly. Our results suggest that a majority label strategy is a viable option. It rendered a higher reliability (48.40 kappa) than the average (47.97) across the three individual runs. This result indicates we can minimize the risk of using an under-performing example order when a validation set is unavailable simply by running the LLM multiple times with different orders and then picking the majority label for each example returned by most runs. However, once again, the recommended approach is manually annotating a validation set and evaluating different random orders to maximize performance.